40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 31 21:12:05 2013, manasa, Update, General, Pump down called off 
    Reply  Thu Aug 1 16:03:52 2013, Steve, Update, General, reopen BS & IOO IOOnorth.jpg
    Reply  Thu Aug 1 22:49:31 2013, manasa, Update, General, Green status after PR3 flipping investigated 
       Reply  Fri Aug 2 17:50:10 2013, Jenne, Update, General, PR3 wedge angle adjusted 
          Reply  Fri Aug 2 22:51:10 2013, Jenne, Update, General, All vent tasks complete, just need oplev check 
             Reply  Mon Aug 5 18:02:34 2013, manasa, Update, General, Close up list checked 
             Reply  Tue Aug 6 07:52:58 2013, manasa, Update, General, Ready to pump down MCspot_0806.pngOSEM_0806.pngOL_0806.pngQPD_0806.png
Message ID: 8960     Entry time: Fri Aug 2 17:50:10 2013     In reply to: 8958     Reply to this: 8962
Author: Jenne 
Type: Update 
Category: General 
Subject: PR3 wedge angle adjusted 

[Jenne, Manasa, Koji]

Earlier today, we locked and aligned both the X and Y arms. 

I then went into the BS chamber, put on the BS' aperture, and put an aperture along the AS path.  (We had Michelson fringes, so I centered the aperture around the fringes.  I used one of the brass ruler things that we use to center the beam on ITMs and ETMs, on a riser.  I put this aperture at the edge of the BS table, after the AS beam is launched toward the OMC chamber.  The idea was to replace PR3 such that I could get the beam back through the BS aperture, and the brass ruler aperture, in hopes that we would see arm flashes, and not have to open the ITMY and ETMY heavy doors.)

I set references on the table so that I could put PR3 back in its original position, then removed PR3 from the chamber.

Steve set up a HeNe for me, that we pointed through the optic.  The ghost beam was very high, indicating (as expected) that the wedge was not perfectly horizontal.

I took the suspension off of the cage and laid it down, as I have in the past. 

I removed the optic from the suspension, to try to figure out which was the fat vs. skinny side.  I noticed that there are very faint marks on the actual fat and skinny sides of the optic.  (Mpral - for the LaserOptik mirrors, look for the faint lines that are the full width of the barrel, not the placement of the arrow which marks the HR side).  I put the optic back in (HR side toward the back, fat side on the left (as you look at the face of the optic), which is consistent with the picture in the Optical Layout page of the Wiki, near the bottom.) the optic holder ring.

I put the suspension back on the cage, and saw that the HeNe's ghost beam was now nearly horizontal relative to the straight-through beam.  Excellent.  Also, the pitch balancing didn't seem to change noticably, which I determined was within "poking" distance of where we need it to be.

I put PR3 back onto the BS table, and adjusted it around until I got the beam through both the BS aperture, and the one on the AS path.  As usual, this took quite a while, but as soon as I got through both of those apertures (really at the same place, not close to being through them, but as close as I could tell by eye - this is what took forever), Koji and Manasa saw flashes in the Yarm!  Yay! 

Since I had to move PR3 in angle a tiny bit, I reset the references, then dogged down PR3.  We still had flashes, this time in both arms, so we closed up the light doors.

We have now locked and aligned both arms in IR after the adjustment of PR3, and see both arms' green at 01 or 02. We are about to start checking the green positioning on the periscopes.  We will also need to check the AS path, as well as IPPOS and IPANG before we close up.  We see REFL on the camera.

Separately - Manasa remembered that 2 clean things were dropped yesterday - a screw, and an allen key.  Since they're both Clean, we're not too worried, although she thinks a long-armed person may be able to reach the allen key.

ELOG V3.1.3-